# Learning Topology-Driven Multi-Subspace Fusion for Grassmannian Deep Networks

### v1

@@ Introduction
In recent years, unitary subspace modelling on the Grassmannian has achieved great success in tasks which require extended discriminative capacity, such as activity recognition (cherian2017generalized), emotion recognition (liu2014combining), face verification (huang2015projection), and classification of time-series data in brain-computer interfaces (gao2022domain). In the above tasks, Grassmannian typically model the input data with a single and static orthogonal subspace (huang2018building; wang2024grassmannian). However, A fixed single-subspace assumption fails to capture local geometric variations and multi-modal distribution, limiting the expressiveness and adaptability of the model in structurally diverse or heterogeneous task scenarios (hamm2008grassmann). To overcome this issue, we introduce an adaptive multi-subspace representation that dynamically adjusts multiple subspaces to accurately capture task-specific characteristics and the harmonious distribution among subspaces. The convergence of obtaining adaptive subspaces is analysed from a topological perspective (munkres2000topology), satisfying specific demands across different tasks.
@@
In the Euclidean space, the promising progress of deep learning in the recent decade is attributed to multi-channel interactions and non-linear activations (krizhevsky2012imagenet). In particular, the seminal work LeNet-5 (lecun2002gradient) extracts multi-channel feature maps using stacked convolution blocks, thus obtaining extended representative capacity via channel interaction. Admitting the natural representation power of linear subspace, the potential of multi-subspace interactions has been overlooked in the community. To reveal the power of multi-subspace interaction for the Grassmannian, there are two major issues, i.e., how to perform subspace interaction and how to formulate its deep architecture.
For the first issue, we design a multi-subspace interaction block in a geometric manner, i.e., several intermediate representations are generated by a unified mapping and then fused through alignment, facilitating effective interaction among different Grassmannian subspaces. For the second issue, we leverage an intrinsic geometric approach to ensure that the fusion process occurs at accurate positions on the manifold (tao2024alignmif), which enables the design of our multi-subspace interaction block in a stackable form, allowing for flexible capacity expansion. To the best of our knowledge, this is the first work that introduces deep interactions of Grassmannian subspaces in Riemannian neural networks, as shown in Figure 1 (c) and (f), these deep interactions capture the dynamic relationships between multiple subspaces, effectively modelling their geometric structures, semantic features, and mutual influences. Euclidean deep networks naturally implement backpropagation, given that the gradient superposition property satisfies the axioms of linear spaces (fei2025survey). However, parameter optimisation on a Riemannian manifold must be strictly restricted to its tangent space  (smith2014optimization). To this end, we develop a topology-driven framework for adaptive multi-subspace construction on the Grassmannian. In summary, our main contributions are the following:
@@
Optimisation Strategies: In order to effectively train the proposed GMSF-Net, we restructure the loss function by combining classification cross-entropy loss (hinton1995wake) and mutual information loss, effectively capturing diverse geometric subspaces. The total loss function for the proposed method is as follows:
@@
where, $\mathcal{L}_{total}$ is the total loss, $\mathcal{L}_{CE}$ is the cross-entropy loss, $\mathcal{L}_{C}$ is the mutual information loss, and $\lambda$ is a hyperparameter that balances classification accuracy with structural preservation. The details of these two loss functions (Appendix D.1) and their backpropagation derivations (Appendix D.2) are provided in Appendix D.

### v2

@@ Abstract
Codehttps://github.com/Xua-Yu/GMSF-Net
Extended versionhttps://arxiv.org/abs/2511.08628
@@ Introduction
In recent years, unitary subspace modelling on the Grassmannian has achieved great success in tasks which require extended discriminative capacity, such as activity recognition (cherian2017generalized), emotion recognition (liu2014combining), face verification (huang2015projection), and classification of time-series data in brain-computer interfaces (gao2022domain). In the above tasks, Grassmannian typically model the input data with a single and static orthogonal subspace (huang2018building; wang2024grassmannian; xu2023learning). However, a fixed single-subspace assumption fails to capture local geometric variations and multi-modal distribution, limiting the expressiveness and adaptability of the model in structurally diverse or heterogeneous task scenarios (hamm2008grassmann). To overcome this issue, we introduce an adaptive multi-subspace representation that dynamically adjusts multiple subspaces to accurately capture task-specific characteristics and the harmonious distribution among subspaces. The convergence of obtaining adaptive subspaces is analysed from a topological perspective (munkres2000topology), satisfying specific demands across different tasks.
@@
In the Euclidean space, the promising progress of deep learning in the recent decade is attributed to multi-channel interactions and non-linear activations (krizhevsky2012imagenet). In particular, the seminal work LeNet-5 (lecun2002gradient) extracts multi-channel feature maps using stacked convolution blocks, thus obtaining extended representative capacity via channel interaction. Admitting the natural representation power of linear subspace (xu2018non), the potential of multi-subspace interactions has been overlooked in the community. To reveal the power of multi-subspace interaction for the Grassmannian, there are two major issues, i.e., how to perform subspace interaction and how to formulate its deep architecture.
For the first issue, we design a multi-subspace interaction block in a geometric manner, i.e., several intermediate representations are generated by a unified mapping and then fused through alignment, facilitating effective interaction among different Grassmannian subspaces. For the second issue, we leverage an intrinsic geometric approach to ensure that the fusion process occurs at accurate positions on the manifold (tao2024alignmif), which enables the design of our multi-subspace interaction block in a stackable form, allowing for flexible capacity expansion (song2025refinefuse). To the best of our knowledge, this is the first work that introduces deep interactions of Grassmannian subspaces in Riemannian neural networks, as shown in Figure 1 (c) and (f), these deep interactions capture the dynamic relationships between multiple subspaces, effectively modelling their geometric structures, semantic features, and mutual influences. Euclidean deep networks naturally implement backpropagation, given that the gradient superposition property satisfies the axioms of linear spaces (fei2025survey). However, parameter optimisation on a Riemannian manifold must be strictly restricted to its tangent space  (smith2014optimization). To this end, we develop a topology-driven framework for adaptive multi-subspace construction on the Grassmannian. In summary, our main contributions are the following:
@@
Optimisation Strategies: In order to effectively train the proposed GMSF-Net, we restructure the loss function by combining classification cross-entropy loss (hinton1995wake) and MI-inspired regularizer loss, effectively capturing diverse geometric subspaces. The total loss function for the proposed method is as follows:
@@
where, $\mathcal{L}_{total}$ is the total loss, $\mathcal{L}_{CE}$ is the cross-entropy loss, $\mathcal{L}_{C}$ is the MI-inspired regularizer loss, and $\lambda$ is a hyperparameter that balances classification accuracy with structural preservation. The details of these two loss functions (Appendix D.1) and their backpropagation derivations (Appendix D.2) are provided in Appendix D.